What is P=3qr solved for q?
A. 3rp=q *
B. P/3r=q
C. 3r/P=q
DP/3q=r

Answer by Cromlix(4381)

Hi there,
P = 3qr
Divide both sides by 3r
P/3r = 3qr/3r
P/3r = q (B)
Hope this helps :-)